OpenStudy (anonymous):

what value of C makes the polynomial below a perfect square trinomial? 4x^2-36x+c

OpenStudy (anonymous):

now we are going the other way what is 4 the square of?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

careful not "what is 4 squared" but rather "what number squared is 4?"

OpenStudy (anonymous):

ohhhhh 2

OpenStudy (anonymous):

ok good now recall that \[a^2+2ab+b^2=(a+b)^2\] we have \(a=2, 2ab=2\times 2\times b=4b=-36\) so what is \(b\) ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

well actually -9 but no matter if \(b=9\) then what is \(b^2\) ?

OpenStudy (anonymous):

-81

OpenStudy (anonymous):

always plus for these it is \(81\) and that is your answer

OpenStudy (anonymous):

thanks bro agian xD

OpenStudy (anonymous):

\[4x^2-36x+81=(2x-9)^2\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

(again)